Use the method of Frobenius to obtain linearly independent series solutions about r = 0. 1.0y + 1ry + (22 – 1)y=0. Use an
Answer #1

n-0 n= 0 no 18x® y +ay + (-17-0 0 Frobenius series solution about a = xo is yle) = Cn (2 - 0)*** Here to = 0 :.(x) = @ Chan(ntr) Cn anty - Z (nto) (n+8-1) Co anty + no no n=o Co znova - Centre - 0 3 [car)\nex 1) +18) 16 entry nants n=0 + 8 in antr+. . In 1st sum, - put coefficient for n=0 equals to zero. .. (0+8-1) (0 +Y+1) Co = 0 Take Co to :(8-1)(8+1) = 0 > 8=d.at :. In=o .☺ becomes, nto co ntriga (ntr-1)(n+8+1) Cn Xto + E Cn achtsta=0 n=2 40 1. Now, in second sum of B ② replace n by n-a Iwe-> (n+8-1) (1 +8+1) Cn + Cn-g = 0 in=2,3,--. =) (n = in =2,3, --- - Cn-q intr-1)(n+8+)) Larger root in Vg at in = 2, 3, -- -k=2 k2 grootste K=3 K-3 (z = =0 (as C, =0) 3(3+2) k=5 Cg = -6 g = 0 (as (3=0) 5x 7 I and so on . and so on. as yra) = 3 Cn xhYola) = a (60+0.37 (+4)2° +0.X? t ap 49 70 x3 +.-) :: 4 (2) = 26061 70-34 g 70.22 + test 70.3 + -) : Co(2 +0.27-32 70.49 te t